# Hubbard and Heisenberg models on hyperbolic lattices - Metal-insulator transitions, global antiferromagnetism and enhanced boundary fluctuations

### v1

@@ Appendix B Spin wave approximation
$\frac{d}{dt}\hat{\bm{b}}=i\begin{pmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{pmatrix}\hat{\bm{b}}$ (24)
@@
$\hat{\bm{b}}^{\dagger}$ $=$ $(\hat{b}_{\bm{i}=1}^{\dagger},\dots,\hat{b}_{\bm{i}=N}^{\dagger},\hat{b}_{\bm{i}=1}^{\phantom{\dagger}},\dots,\hat{b}_{\bm{i}=N}^{\phantom{\dagger}})\,,$
@@
In order to diagonalize the Hamiltonian of Eq. (23) we perform a Bogoliubov transformation
@@
To ensure, that the new boson operators $\hat{d}_{\bm{i}}^{\dagger}$, that we obtain from $\hat{b}_{\bm{i}}^{\dagger}=\bar{f}_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\phantom{\dagger}}+g_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\dagger}$, still fulfill the canonical commutation relations, we need to impose a condition on the entries of $f$ and $g$
@@
We now define a scalar product $\langle\bm{x},\bm{y}\rangle_{B}=\bm{x}^{\dagger}\left(\begin{smallmatrix}-1&0\\ 0&1\\ \end{smallmatrix}\right)\bm{y}$. The matrix $M=\left(\begin{smallmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{smallmatrix}\right)$ is hermitian with respect to this scalar product $\langle M\bm{x},\bm{y}\rangle_{B}=\langle\bm{x},M\bm{y}\rangle_{B}$ and the Bogoliubov transformation $U=\left(\begin{smallmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{smallmatrix}\right)$ is unitary $\langle U\bm{x},U\bm{y}\rangle_{B}=\langle\bm{x},\bm{y}\rangle_{B}$. After diagonalizing the matrix $M$, the condition on the Bogoliubov transformation of Eq. (27) can be satisfied by ensuring, that all eigenvectors of $M$ are orthonormal with respect to the scalar product $\langle\cdot,\cdot\rangle_{B}$. If the norm of an eigenvector $\langle\bm{x}_{n},\bm{x}_{n}\rangle_{B}>0$ is nonzero, it can also be shown, that the corresponding eigenvalue is real. In order to avoid eigenvectors with a vanishing $B$ norm we add an infinitesimal staggered magnetic field $\hat{H}_{\epsilon}=-\epsilon\sum_{\bm{i}}(-1)^{\bm{i}}\hat{S}_{\bm{i}}^{z}$. After the calculation, limits have to be taken carefully, by first taking the thermodynamic limit and then the limit of $\epsilon\rightarrow 0$. The site-dependent correction to the classical AFM state can be calculated by the mean boson occupation number at $T=0$
@@
Appendix C Spin susceptibility
@@
Without loss of generality we can consider the staggered spin susceptibility in $x$-direction
@@
As apparent, the partial particle-hole transformation maps the staggered susceptibility in the particle-hole channel to the uniform one in the particle-particle channel. The latter is nothing but the Cooper instability, that we now show to be present on any graph with finite DOS at the Fermi energy. To do so, we introduce a canonical transformation $U$, that diagonalizes the Hamiltonian
$\hat{H}_{t}=\sum_{\bm{i},\bm{j},\sigma}h_{\bm{i},\bm{j}}\hat{\gamma}_{\bm{i},\sigma}^{\dagger}\hat{\gamma}_{\bm{j},\sigma}^{\phantom{\dagger}}=\sum_{x}\lambda_{x}\hat{\eta}_{x,\sigma}^{\dagger}\hat{\eta}_{x,\sigma}^{\phantom{\dagger}}$ (31)
@@
$=$ $\frac{1}{N_{s}}\sum_{x,y}\frac{f_{x,y}}{\lambda_{x}+\lambda_{y}}\left[\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}+\tanh{\left(\frac{\beta}{2}\lambda_{y}\right)}\right]$
with $f_{x,y}=\sum_{\bm{i},\bm{j}}(U^{\dagger})_{x,\bm{i}}U_{\bm{j},x}(U^{\dagger})_{y,\bm{i}}U_{\bm{j},y}$. If the DOS is finite at zero energy, the Fermi level here, we can show, that the term with $x=y$ has a logarithmic divergency
$\chi^{x}(Q)$ $=$ $\frac{1}{N_{s}}\sum_{x}\frac{f_{x,x}}{\lambda_{x}}\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}+R$ (33)
@@
where all other terms in the sum with $x\neq y$ are summarized in the term $R$. The last approximation is in the low-temperature limit and $W$ is a high-energy cutoff. In principle terms could appear in $R$, that lead to the cancellation of the logarithmic divergency under certain conditions. But in general the occurrence of an instability towards AFM ordering is generic even in the absence of translational symmetry.
@@ Appendix D Additional data
In the following section we present some additional data.
@@ D.1 Mean-field approximation
Figure 13: (I) Local DOS $\rho(E=0,\bm{i})$ at zero energy in the non-interacting limit $U=0$ at $\delta=0.01$ and (II) magnetization $m_{\bm{i}}^{z}$ as obtained from mean-field approximation for the $\{10,3\}$ lattice and interaction strength $U=0.8t$.
@@
In Sec. IV.2 we showed the spatial distribution of the magnetization $m_{\bm{i}}^{z}$ in the weak-coupling limit for the $\{8,3\}$ lattice with $n_{l}=5$ layers. In the upper panels of Figs. 13-15 we plot the local DOS at zero energy
$\rho(E,\bm{i})=-\frac{1}{\pi N_{s}}\sum_{n}|U_{\bm{i},n}|^{2}\mathrm{Im}G^{R}(n,E)$ (34)
@@
Here we defined a canonical transformation $U$, that diagonalizes the mean-field Hamiltonian (4). They all show a unique pattern of low and high DOS, that is compatible with the $C_{p}$ symmetry of the lattices. The lower panels of Figs. 13-15 show the corresponding local order parameter in real-space, where again local ferromagnetic and AFM moments emerge in the different sectors of the lattice. But in sum of all lattice sites only the AFM moment remains.
D.2 Spin wave approximation
The results for the site-dependent correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to the classical Néel state obtained from the spin wave approximation for the $\{10,3\}$ and $\{8,3\}$ lattice in Figs. 16(I) and 16(II) are similar to the results for the $\{8,3\}$ lattice described in the main text [Fig. 8]. For the $\{8,8\}$ lattice the results differ slightly due to the lattice geometry. The correction on many sites in the outermost layer is smaller than for the other two lattices, since the distance from one site on the edge to another site on the edge is in general smaller. This is a consequence of the high coordination number $q=8$, since sites in the bulk can have direct connections to the edge. The sites on the edge also form short one-dimensional chains, but larger effective interactions are present in comparison to the other two considered lattice geometries.
Figure 16: Local correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to classical Néel state as obtained from spin wave approximation for the (I) $\{10,3\}$, (II) $\{8,3\}$ and (III) $\{8,8\}$ lattice and varying system size $n_{l}$ at $\epsilon=0.01$.
@@ D.3 Quantum Monte Carlo
Figure 17: AFM order parameter for the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice obtained by QMC calculations at $\beta t=100$ (filled symbols) and $\beta t=25$ (unfilled symbols).
In order to supplement the picture given in Figs. 9 and 11, we present the order parameter $m_{AFM}^{z}$ as a function of $U$ in Fig. 17 for two different temperatures $\beta t=25$ and $\beta t=100$. We see that the $\{8,8\}$ lattice exhibits stronger temperature effects than the other lattices. This lattice shows a strong sensitivity to temperature already in the weak coupling regime, as can be seen in Fig. 17(c). As discussed before, this effect is also present in the results from a mean-field analysis [Fig. 4(a)].

### v2

@@ Appendix A Spin-wave approximation
$\frac{d}{dt}\hat{\bm{b}}=i\begin{pmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{pmatrix}\hat{\bm{b}}\,,$ (27)
@@
$\hat{\bm{b}}^{\dagger}$ $=$ $(\hat{b}_{\bm{i}=1}^{\dagger},\dots,\hat{b}_{\bm{i}=N_{s}}^{\dagger},\hat{b}_{\bm{i}=1}^{\phantom{\dagger}},\dots,\hat{b}_{\bm{i}=N_{s}}^{\phantom{\dagger}})\,,$
@@
To diagonalize the Hamiltonian of Eq. (26) we perform a Bogoliubov transformation,
@@
To ensure that the new boson operators $\hat{d}_{\bm{i}}^{\dagger}$, which we obtain from $\hat{b}_{\bm{i}}^{\dagger}=\bar{f}_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\phantom{\dagger}}+g_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\dagger}$, still fulfill the canonical commutation relations, we need to impose a condition on the entries of $f$ and $g$:
@@
We now define a scalar product $\langle\bm{x},\bm{y}\rangle_{B}=\bm{x}^{\dagger}\left(\begin{smallmatrix}-1&0\\ 0&1\\ \end{smallmatrix}\right)\bm{y}$. The matrix $M=\left(\begin{smallmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{smallmatrix}\right)$ is hermitian with respect to this scalar product $\langle M\bm{x},\bm{y}\rangle_{B}=\langle\bm{x},M\bm{y}\rangle_{B}$ and the Bogoliubov transformation $U=\left(\begin{smallmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{smallmatrix}\right)$ is unitary $\langle U\bm{x},U\bm{y}\rangle_{B}=\langle\bm{x},\bm{y}\rangle_{B}$. After diagonalizing the matrix $M$, the condition on the Bogoliubov transformation of Eq. (30) can be satisfied by ensuring, that all eigenvectors of $M$ are orthonormal with respect to the scalar product $\langle\cdot,\cdot\rangle_{B}$. If the norm of an eigenvector $\langle\bm{x}_{n},\bm{x}_{n}\rangle_{B}>0$ is nonzero, then it can also be shown that the corresponding eigenvalue is real. To avoid eigenvectors with a vanishing $B$ norm we add an infinitesimal staggered magnetic field $\hat{H}_{\epsilon}=-\epsilon\sum_{\bm{i}}(-1)^{\bm{i}}\hat{S}_{\bm{i}}^{z}$. After the calculation, limits have to be taken carefully, by first taking the thermodynamic limit and then the limit of $\epsilon\rightarrow 0$. The site-dependent correction to the classical AFM state can be calculated by the mean boson occupation number at $T=0$:
@@
Appendix B Staggered spin susceptibility
@@
Without loss of generality we can consider the staggered spin susceptibility in the $x$ direction:
@@
As apparent, the partial particle-hole transformation maps the staggered susceptibility in the particle-hole channel to the uniform one in the particle-particle channel. The latter is nothing but the Cooper instability, which we now show to be present on any graph with finite DOS at the Fermi energy. To do so, we introduce a canonical transformation $U$, that diagonalizes the Hamiltonian
$\hat{H}_{t}=\sum_{\bm{i},\bm{j},\sigma}h_{\bm{i},\bm{j}}\hat{\gamma}_{\bm{i},\sigma}^{\dagger}\hat{\gamma}_{\bm{j},\sigma}^{\phantom{\dagger}}=\sum_{x}\lambda_{x}\hat{\eta}_{x,\sigma}^{\dagger}\hat{\eta}_{x,\sigma}^{\phantom{\dagger}}\,,$ (34)
@@
$=$ $\frac{1}{N_{s}}\sum_{x,y}\frac{f_{x,y}}{\lambda_{x}+\lambda_{y}}\left[\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}+\tanh{\left(\frac{\beta}{2}\lambda_{y}\right)}\right]\,,$
with $f_{x,y}=\sum_{\bm{i},\bm{j}}(U^{\dagger})_{x,\bm{i}}U_{\bm{j},x}(U^{\dagger})_{y,\bm{i}}U_{\bm{j},y}$. For our specific case, the hopping Hamiltonian is symmetric such that $U$ is an orthogonal matrix. As a consequence, $f_{x,y}=\delta_{x,y}$. Hence,
$\chi^{x}(Q)$ $=$ $\frac{1}{N_{s}}\sum_{x}\frac{1}{\lambda_{x}}\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}$
[…]
where the DOS is given by Eq. (13). The last approximation is in the low-temperature limit and $W$ is a high-energy cutoff. Provided that the DOS is finite at the Fermi energy, an instability toward AFM ordering is generic even in the absence of translational symmetry.
Appendix C Instability of the bulk
To study the stability of the bulk toward AFM ordering, we place an AFM perturbation $\hat{H}_{I}$ on the inner ring of the lattice
$\hat{H}$ $=$ $\hat{H}_{t}+\hat{H}_{I}\,,$ (37)
[…]
$\hat{H}_{I}$ $=$ $\sum_{i,\sigma}m\sigma(-1)^{i}\hat{c}^{\dagger}_{i,1,\sigma}\hat{c}^{\phantom{\dagger}}_{i,1,\sigma}=\sum_{k,\sigma}m\sigma\hat{c}^{\dagger}_{k,1,\sigma}\hat{c}^{\phantom{\dagger}}_{k+Q,1,\sigma}\,.$
In Eq. (37), we used the notation as introduced in Eq. (IV.2.2), where we defined a unit cell $i$ with the help of the $C_{p}$ symmetry of the lattice. The orbital $n=1$ refers to the site on the inner layer included in the unit cell $i$. Equation (37) is equivalent to Eq. (IV.2.2) by setting $M_{1,1}=m$ and $M_{n,n^{\prime}}=0$ otherwise.
The susceptibility with respect to the perturbation $\hat{H}_{I}$ can be written as
@@ Appendix C Instability of the bulk
[…]
×$\times A_{1,1}(k,\omega)A_{1,1}(k+Q,\omega^{\prime})\,.$
Here, we defined the Green function $G_{1,1}(k,\tau)=-\langle\mathcal{T}\hat{c}_{k,1}^{\phantom{\dagger}}(\tau)\hat{c}_{k,1}^{\dagger}\rangle_{0}$ with the imaginary-time ordering operator $\mathcal{T}$ and the expectation value $\langle\dots\rangle_{0}$ with respect to $\hat{H}_{t}$. The spectral function $A_{1,1}(k,\omega)=-\frac{1}{\pi}\mathrm{Im}G_{1,1}(k,i\omega_{m}\rightarrow\omega+i\delta)$ is defined via analytical continuation of the Green function with an infinitesimal $\delta$ and the fermionic Matsubara frequencies $\omega_{m}$. Using particle-hole symmetry, $A_{1,1}(k,\omega)=A_{1,1}(k+Q,-\omega)$, we can further simplify the last equation to obtain:
[…]
×$\times\sum_{k}A_{1,1}(k,\omega)A_{1,1}(k,\omega^{\prime})\,.$
Since $A_{1,1}(k,\omega)$ takes a finite, constant value at the Fermi energy [27] we can replace the $k$-sum over the spectral functions by a finite number. Carrying out the integration leads to a finite value of $\chi(Q,i\Omega_{m}=0)$ such that no bulk magnetic instability is present.
@@ Appendix D Additional data
In the following Appendix we present some additional data.
@@ D.1 Mean-field approximation
Figure 14: (I) Local DOS $\rho(E=0,\bm{i})$ at zero energy in the noninteracting limit $U=0$ at $\delta=0.01$ and (II) magnetization $m_{\bm{i}}^{z}$ as obtained from mean-field approximation for the $\{10,3\}$ lattice and interaction strength $U=0.8t$.
@@
In Sec. IV.2 we showed the spatial distribution of the magnetization $m_{\bm{i}}^{z}$ in the weak-coupling limit for the $\{8,3\}$ lattice with $n_{l}=5$ layers. In the upper panels of Figs. 1416 we plot the local DOS at zero energy,
$\rho(E,\bm{i})=-\frac{1}{\pi N_{s}}\sum_{n}|U_{\bm{i},n}|^{2}\mathrm{Im}G^{R}(n,E)\,,$ (40)
@@
Here we defined a canonical transformation $U$, that diagonalizes the mean-field Hamiltonian (4). They all show a unique pattern of low and high DOS, which is compatible with the $C_{p}$ symmetry of the lattices. The lower panels of Figs. 1416 show the corresponding local order parameter in real space, where again local ferromagnetic and AFM moments emerge in the different sectors of the lattice. But in sum of all lattice sites only the AFM moment remains.
D.2 Spin-wave approximation
The results for the site-dependent correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to the classical Néel state obtained from the spin-wave approximation for the $\{10,3\}$ and $\{8,3\}$ lattice in Figs. 17(I) and 17(II) are similar to the results for the $\{8,3\}$ lattice described in the main text [Fig. 9]. For the $\{8,8\}$ lattice the results differ slightly due to the lattice geometry. The correction on many sites in the outermost layer is smaller than for the other two lattices, since the distance from one site on the edge to another site on the edge is in general smaller. This is a consequence of the high coordination number $q=8$, since sites in the bulk can have direct connections to the edge. The sites on the edge also form short one-dimensional chains, but larger effective interactions are present in comparison to the other two considered lattice geometries.
Figure 17: Local correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to classical Néel state as obtained from spin-wave approximation for the (I) $\{10,3\}$, (II) $\{8,3\}$, and (III) $\{8,8\}$ lattice and varying system size $n_{l}$ at $\epsilon=0.01$.
@@ D.3 Quantum Monte Carlo
Figure 18: AFM order parameter for the $\{10,3\}$, $\{8,3\}$, and $\{8,8\}$ lattice obtained by QMC calculations at $\beta t=100$ (filled symbols) and $\beta t=25$ (unfilled symbols).
To supplement the picture given in Figs. 10 and 12, we present the order parameter $m_{\mathrm{AFM}}^{z}$ as a function of $U$ in Fig. 18 for two different temperatures $\beta t=25$ and $\beta t=100$. We see that the $\{8,8\}$ lattice exhibits stronger temperature effects than the other lattices. This lattice shows a strong sensitivity to temperature already in the weak coupling regime, as can be seen in Fig. 18(c). As discussed before, this effect is also present in the results from a mean-field analysis [Fig. 5(a)].